### v4

# Algebraic similarity

Abstract.
Detecting and exploiting similarities between seemingly distant objects is without doubt an important human ability. This paper develops from the ground up an abstract algebraic and qualitative justification-based notion of similarity based on the observation that sets of generalizations encode important properties of elements. We show that similarity defined in this way has appealing mathematical properties. As we construct our notion of similarity from first principles using only elementary concepts of universal algebra, to convince the reader of its plausibility, we show that it can be naturally embedded into first-order logic via model-theoretic types.
1. Introduction
The purpose of this paper is to introduce from first principles an abstract algebraic and qualitative justification-based notion of similarity based on sets of generalizations as motivated by the following observations. We say that a term $s(x_{1},\ldots,x_{n})$ is a generalization of a natural number $a\in\mathbb{N}$ iff there is some sequence of natural numbers $o_{1},\ldots,o_{n}$ such that $a=s(o_{1},\ldots,o_{n})$. For example, the term $2x$ is a generalization of $6$ since $6=2\cdot 3$. More generally speaking, the term $2x$ is a generalization of every even number, that is,
$2x\text{ is a generalization of $a\in\mathbb{N}$}\quad\Leftrightarrow\quad a\text{ is even}.$
We also have
$x^{2}\text{ is a generalization of $a\in\mathbb{N}$}\quad\Leftrightarrow\quad a\text{ is a square number}.$
These examples indicate that generalizations encode important properties of the generalized elements which motivates the following definition of generalization-based similarity. Let $\mathfrak{A}$ and $\mathfrak{B}$ be algebras with identical underlying language of algebras $L$, and let $a$ and $b$ be elements of the universes of $\mathfrak{A}$ and $\mathfrak{B}$, respectively. We denote the set of all generalizations of $a$ and $b$ by $\uparrow_{\mathfrak{A}}a$ and $\uparrow_{\mathfrak{B}}b$, respectively. Since these sets of generalizations encode properties of $a$ and $b$ by the above observations, the intersection
$a\uparrow_{\mathfrak{AB}}b:=(\uparrow_{\mathfrak{A}}a)\cap(\uparrow_{\mathfrak{B}}b)$
contains the joint properties of $a$ and $b$ (since $\mathfrak{A}$ and $\mathfrak{B}$ have the same underlying language, the terms in both sets contain symbols from the same alphabet). We now say that $a$ and $b$ are similar iff this set of joint properties $a\uparrow_{\mathfrak{AB}}b$ is $\subseteq$-maximal with respect to $a$ and $b$ (see Definition 3).
It turns out that similarity as defined in this paper has appealing mathematical properties:
The First 19 and Second Isomorphism Theorem 21 show that similarity is compatible with bijective structure-preserving mappings as desired. Example 20, on the other hand, shows that similarity is in general not compatible with homomorphisms, justified by a simple counterexample.
In §6, we introduce the $(k,\ell)$-fragments of the framework consisting of restricted generalizations and we briefly analyze the monolinear fragment $(1,1)$ — consisting only of generalizations with exactly one occurrence of a single variable — in the domain of sets and numbers.
In §8, we establish a tight connection between similarity in finite algebras and finite tree automata.
As we construct similarity from first principles using only elementary concepts of universal algebra, to convince the reader of its plausibility we need to validate it either empirically or — what we prefer here — theoretically by showing that it fits naturally into the overall mathematical landscape. For this, we will show in §9 that the purely algebraic notion of similarity can be naturally embedded into first-order logic via model-theoretic types. More precisely, we show that sets of generalizations are in one-to-one correspondence with so-called “g-formulas” and “g-types”, which is appealing as types play a fundamental role in model theory and showing that our notion — which is primarily motivated by simple examples — has a natural logical interpretation, provides strong evidence for its suitability.
We do not claim to obtain any deep results — the purpose of this paper is rather to introduce the framework of algebraic similarity and to initiate its study.
2. Preliminaries
Let $\mathbb{N}_{k}:=\{k,k+1,\ldots\}$ be the natural numbers starting at some $k\geq 0$.
We assume the reader to be fluent in basic universal algebra as it is presented for example in ? (?, §II).
A language $L$ of algebras is a set of function symbols [^1] We omit constant symbols as we identify them with 0-ary function symbols. together with a rank function $r:L\to\mathbb{N}$, and a denumerable set $X$ of variables distinct from $L$. Terms are formed as usual from variables in $X$ and function symbols in $L$ and we denote the set of all such $L$-terms by $T_{L}(X)$. The rank of a term $s$ is given by the number of variables occurring in $s$ and is denoted by $r(s)$.
An $L$-algebra $\mathfrak{A}$ consists of a non-empty set $A$, the universe of $\mathfrak{A}$, and for each function symbol $f\in L$, a function $f^{\mathfrak{A}}:A^{r(f)}\to A$, the functions of $\mathfrak{A}$ (the distinguished elements of $\mathfrak{A}$ are the 0-ary functions). Every term $s$ induces a function $s^{\mathfrak{A}}$ on $\mathfrak{A}$ in the usual way.
A homomorphism is a mapping $H:\mathfrak{A\to B}$ between two $L$-algebras satisfying, for each function symbol $f\in L$ and elements $a_{1},\ldots,a_{r(f)}\in A$,
$Hf^{\mathfrak{A}}(a_{1},\ldots,a_{r(f)})=f^{\mathfrak{B}}(Ha_{1},\ldots,Ha_{r(f)}).$
An isomorphism is a bijective homomorphism.
3. Algebraic similarity
This is the main section of the paper. Here, we introduce an abstract algebraic and qualitative notion of similarity based on the observation that sets of generalizations contain important information about elements (see the discussion in §1).
Let $\mathfrak{A}$ and $\mathfrak{B}$ be $L$-algebras over some joint (ranked) language of algebras $L$.
Definition 1.
Define the set of generalizations of an element $a\in A$ by
$\uparrow_{\mathfrak{A}}a:=\left\{s\in T_{L}(X)\;\middle|\;a=s^{\mathfrak{A}}\mathbf{o},\text{ for some $\mathbf{o}\in A^{r(s)}$}\right\},$
extended to elements $a\in A$ and $b\in B$ by
$a\uparrow_{\mathfrak{AB}}b:=(\uparrow_{\mathfrak{A}}a)\cap(\uparrow_{\mathfrak{B}}b).$
Example 2.
In the natural numbers with multiplication $\mathfrak{M}=(\mathbb{N},\cdot,\mathbb{N})$, [^2] The second $\mathbb{N}$ in $(\mathbb{N},\cdot,\mathbb{N})$ indicates that every natural number is a distinguished element and can thus be used as a constant (symbol) to form terms and generalizations. evenness is captured via
$a\in\mathbb{N}\text{ is even }\quad\Leftrightarrow\quad 2x\in\ \uparrow_{\mathfrak{M}}a,$
and primality via
$a\in\mathbb{N}\text{ is prime }\quad\Leftrightarrow\quad\uparrow_{\mathfrak{M}}a=\{x,ax\}.$
The above example shows how generalizations encode properties of the generalized elements. We can thus interpret the set of all generalizations of an element as the set of all properties expressible within the surrounding algebra — the set of shared properties of two elements $a\in A$ and $b\in B$ is therefore captured by $a\uparrow_{\mathfrak{AB}}b$.
The above discussions motivate the following algebraic definition of similarity via generalizations:
Definition 3 (Similarity).
We define the similarity relation in two steps:
  1. (1)
    A generalization is trivial in $\mathfrak{AB}$ iff it generalizes all elements in $A$ and $B$.
    Now we say that $a\lesssim b$ holds in $\mathfrak{AB}$ — in symbols,
    $a\lesssim_{\mathfrak{AB}}b,$
    iff
    1. (a)
      either $(\uparrow_{\mathfrak{A}}a)\cup(\uparrow_{\mathfrak{B}}b)$ consists only of trivial generalizations; or
    2. (b)
[…]
      $\emptyset\subsetneq a\uparrow_{\mathfrak{AB}}b\subseteq a\uparrow_{\mathfrak{AB}}b^{\prime}\quad\Rightarrow\quad a\uparrow_{\mathfrak{AB}}b^{\prime}\subseteq a\uparrow_{\mathfrak{AB}}b.$
      We abbreviate the above requirement by simply saying that $a\uparrow_{\mathfrak{AB}}b$ is $b$-maximal.
  2. (2)
    Finally, the similarity relation is defined as
    $a\approx_{\mathfrak{AB}}b\quad:\Leftrightarrow\quad a\lesssim_{\mathfrak{AB}}b\quad\text{and}\quad b\lesssim_{\mathfrak{BA}}a,$
    in which case we say that $a$ and $b$ are similar in $\mathfrak{AB}$.
We will always write $\mathfrak{A}$ instead of $(\mathfrak{A,A})$.
Notice that $\uparrow$ is symmetric in the sense that
$a\uparrow_{\mathfrak{AB}}b=b\uparrow_{\mathfrak{BA}}a.$
Of course,
(1) $\uparrow_{\mathfrak{A}}a\subseteq\ \uparrow_{\mathfrak{B}}b\quad\Rightarrow\quad a\lesssim_{\mathfrak{AB}}b.$
and
(2) $\uparrow_{\mathfrak{A}}a=\ \uparrow_{\mathfrak{B}}b\quad\Rightarrow\quad a\approx_{\mathfrak{AB}}b.$
We demonstrate similarity by giving some illustrative examples:
Example 4.
The elements of any algebra of cardinality two are similar as an immediate consequence of the “except for” part of the definition.
Example 5.
In any algebra $(A)$ containing no functions, all elements are similar.
Example 6.
In the algebra ($(\{a,b,\ast\},f)$ given by
{tikzpicture}
we have
$\uparrow a=\ \uparrow b=\left\{f^{n}x\;\middle|\;n\geq 0\right\},$
and therefore
$a\approx b.$
Example 7.
Consider the algebra given by
{tikzpicture}
We compute
$\uparrow a$ =$=\{x\},$
$\uparrow b$ =$=\{x,fx\},$
$\uparrow c$ =$=\left\{f^{n}x\;\middle|\;n\geq 0\right\},$
$=\ \uparrow d,$
$=\ \uparrow e.$
This yields with (1) and (2):
$a<b<c\approx d\approx e.$
Example 8.
[…]
$f0:=0\quad\text{and}\quad fn:=\begin{cases}n-1&n>0,\\ n+1&n<0.\end{cases}$
This can be visualized as
{tikzpicture}
We compute
$\uparrow_{(\mathbb{Z},f)}a=\left\{f^{n}x\;\middle|\;n\geq 0\right\},\quad\text{for all $a\in\mathbb{Z}$}.$
We thus have
$a\approx_{(\mathbb{Z},f)}b,\quad\text{for all $a,b\in\mathbb{Z}$}.$
On the other hand, in the algebra $(\mathbb{N},f)$ of the natural numbers (starting at 00) with successor $fa:=a+1$ given by
{tikzpicture}
we have
[…]
which yields
$a\lesssim_{(\mathbb{N},f)}b\quad\Leftrightarrow\quad a\leq b$
and thus
$a\approx_{(\mathbb{N},f)}b\quad\Leftrightarrow\quad a=b.$
Example 9.
In the mulitplicative arithmetical algebra $\mathfrak{M}:=(\mathbb{N}_{2},\cdot,\mathbb{N}_{2})$, where $\mathbb{N}_{2}:=\{2,3,\ldots\}$ are the natural numbers starting at 2, the following counterexample shows that $\lesssim$ is not compatible with the divisibility relation as we are going to prove
(3) $4\text{ divides }8\quad\text{whereas}\quad 4\not\lesssim 8$
We have
$\uparrow 4$ =$=\{x,xy,2x,x^{2},4\},$
$\uparrow 8$ =$=\{x,xy,xyx,2x,4x,8\}.$
Notice that $x^{2}$ is a generalization whereas it is not a generalization of 8. Now we compute the set of generalizations of the square number 16:
$\uparrow 16=\{16,x,xy,xyx,wxyx,2x,2xy,2xyx,4x,4xy,8x,x^{2}\}.$
We see that
$4\uparrow 8=\{x,xy,2x\}\subsetneq\{x,xy,2x,x^{2}\}=4\uparrow 16$
and thus (3) holds; moreover, it shows
$4\lesssim 16$
Example 10.
Interestingly, for every prime number $p$, in $\mathfrak{M}^{\prime}:=(\mathbb{N}_{2},\cdot)$ not containing any distinguished elements, we have
$\uparrow_{\mathfrak{M}^{\prime}}p=\{x\},$
that is, $p$ has only the trivial generalization $x$. Hence,
(4) $p\approx_{\mathfrak{M}^{\prime}}q,\quad\text{for any prime numbers $p$ and $q$.}$
The situation changes if we consider the algebra $\mathfrak{M}$ from Example 9 containing distinguished elements for any number in $\mathbb{N}_{2}$, since we now have
$\uparrow_{\mathfrak{M}}p=\{x,p\}$
containing the non-trivial generalization $p$, and therefore
$p\lesssim_{\mathfrak{M}}a\quad\Leftrightarrow\quad\text{$p$ divides $a$}.$
Hence (compare to (4)),
$p\approx_{\mathfrak{M}}q\quad\Leftrightarrow\quad p=q,\quad\text{for any prime numbers $p$ and $q$}.$
Example 11.
In the word domain ($(\{a\}^{+},\cdot,\{a\}^{+})$, where $\{a\}^{+}$ is the set of all non-empty words over $a$ and $\cdot$ denotes the concatenation of words, the following counterexample shows that $\lesssim$ is not compatible with the prefix relation as we are going to prove:
(5) $aa\text{ is a prefix of }aaa\quad\text{whereas}\quad aa\not\lesssim aaa.$
We have
$\uparrow aa$ =$=\{x,xy,x^{2},aa\}$
$\uparrow aaa$ =$=\{x,xy,xyz,ax,axy,aax,axa,xaa,xa,aaa\}.$
We see that $x^{2}$ is a generalization of $aa$ which is not a generalization of $aaa$. Now it is easy to verify that $\uparrow aaaa$ contains all generalizations of $aa$ (except for the word $aa$ itself), which directly yields
$aa\uparrow aaa\subsetneq aa\uparrow aaaa,$
which proves (5); moreover, it shows
$aa\lesssim aaaa.$
Since ($(\{a\}^{+},\cdot,\{a\}^{+})$ is isomorphic to $(\mathbb{N}_{1},+,\mathbb{N}_{1})$ with $\mathbb{N}_{1}:=\{1,2,\ldots\}$, where “is a prefix of” translates into “is less than or equal to”, we have derived a result similar in spirit to Example 9 where the multiplicative case is treated.
4. Properties
Computing all generalizations is often difficult which fortunately can be omitted in many cases:
Definition 12.
We call a set $G$ of $L$-terms a characteristic set of generalizations of $a\lesssim b$ in $\mathfrak{AB}$ iff $G$ is a sufficient set of generalizations, that is, iff
  1. (1)
    $G\subseteq a\uparrow_{\mathfrak{AB}}b$, and
  2. (2)
    $G\subseteq a\uparrow_{\mathfrak{AB}}b^{\prime}$ implies $b^{\prime}=b$, for each $b^{\prime}\neq a\in B$.
In case $G=\{s\}$ is a singleton set satisfying both conditions, we call $s$ a characteristic generalization of $a\lesssim b$ in $\mathfrak{AB}$.
Definition 13.
For an $L$-term $s$, define
$\downarrow_{\mathfrak{A}}s:=\left\{s^{\mathfrak{A}}\mathbf{o}\in A\;\middle|\;\mathbf{o}\in A^{r(s)}\right\},$
extended to a set of $L$-terms $G$ by
$\downarrow_{\mathfrak{A}}G:=\bigcap_{s\in G}\downarrow_{\mathfrak{A}}s.$
Fact 14.
We have
$G\subseteq\ \uparrow_{\mathfrak{A}}a\quad\Leftrightarrow\quad a\in\ \downarrow_{\mathfrak{A}}G,$
and $G$ is a characteristic set of generalizations of $a\lesssim b$ in $\mathfrak{AB}$ iff $a\in\ \downarrow_{\mathfrak{A}}G$ and $\downarrow_{\mathfrak{B}}b=\{b\}$.
Example 15.
For any subset $A\subseteq U$ of a given universe $U$, and any element $a\in U$, we have in $(2^{U},\cup,\cap,.^{c})$ the similarity
$A\approx(A\cup\{a\})$
with the characteristic generalization
$s(X):=(A\cup(X\cap\{a\}))$
since
$\downarrow s=\{A,A\cup\{a\}\}$
and see Fact 14.
We now analyze the following basic properties:
[…]
Theorem 16.
The similarity relation is reflexive, symmetric, and in general not transitive.
Proof.
Reflexivity and symmetry hold trivially.
To disprove transitivity, we first consider the three unary algebras
$\mathfrak{A}=(\{a,a^{\prime}\},f,g)\quad\text{and}\quad\mathfrak{B}=(\{b\},f,g)\quad\text{and}\quad\mathfrak{C}=(\{c,c^{\prime}\},f,g)$
given by
{tikzpicture}
Since $b$ is the only element in $\mathfrak{B}$, we must have
$a\lesssim_{\mathfrak{AB}}b.$
Moreover,
$gfx\in b\uparrow_{(\mathfrak{B,C})}c\quad\text{whereas}\quad gfx\not\in b\uparrow_{(\mathfrak{B,C})}c^{\prime}$
shows the $c$-maximality of $b\uparrow_{(\mathfrak{B,C})}c$ which implies
$b\lesssim_{(\mathfrak{B,C})}c.$
On the other hand,
$a\uparrow_{(\mathfrak{A,C})}c^{\prime}=\ \uparrow_{\mathfrak{A}}a$
and
$fgx\in a\uparrow_{(\mathfrak{A,C})}c^{\prime}\quad\text{whereas}\quad fgx\not\in a\uparrow_{(\mathfrak{A,C})}c$
shows
$a\uparrow_{(\mathfrak{A,C})}c\subsetneq a\uparrow_{(\mathfrak{A,C})}c^{\prime}$
which implies
$a\not\lesssim_{(\mathfrak{A,C})}c.$
The proof can easily be adapted to the single algebra $\mathfrak{D}$ consisting of the three algebras $\mathfrak{A,B,C}$ which means that transitivity may fail even in a single algebra. inline]Der Beweis ist noch nicht ganz fertig, oder!? ∎
5. Isomorphism Theorems
It is reasonable to expect isomorphisms — which are bijective structure-preserving mappings between algebras — to be compatible with similarity.
Example 17.
Let $A:=\{a\}$ be the alphabet consisting of the single letter $a$, and let $A^{\ast}$ denote the set of all words over $A$ including the empty word $\varepsilon$. We can identify every sequence $a^{n}=a\ldots a$ ($n$ consecutive $a$’s) with the non-negative integer $n$, for every $n\geq 0$. Therefore, define the isomorphism $H:(\mathbb{N},+)\to(A^{\ast},\cdot)$ via
$H0:=\varepsilon\quad\text{and}\quad Hn:=a^{n},\quad n\geq 1.$
We expect
$n\approx_{((\mathbb{N},+),(A^{\ast},\cdot))}Hn,\quad\text{for all $n\geq 0$.}$
That this is indeed the case is the content of the First Homomorphism Theorem 19 below.
Lemma 18 (Homomorphism Lemma).
For any homomorphism $H:\mathfrak{A\to B}$ and element $a\in A$, we have
$\uparrow_{\mathfrak{A}}a\subseteq\ \uparrow_{\mathfrak{B}}Ha.$
In case $H$ is an isomorphism, we have
$\uparrow_{\mathfrak{A}}a=\ \uparrow_{\mathfrak{B}}Ha.$
Proof.
Since $H$ is a homomorphism, we have
$s\in\ \uparrow_{\mathfrak{A}}a\quad$ $\Leftrightarrow\quad a=s^{\mathfrak{A}}\mathbf{o},\quad\text{for some $\mathbf{o}\in A^{r(s)}$},$
$\Rightarrow\quad Ha=Hs^{\mathfrak{A}}\mathbf{o}=s^{\mathfrak{B}}H\mathbf{o},$
$\Rightarrow\quad s\in\ \uparrow_{\mathfrak{B}}Ha.$
In case $H$ is an isomorphism, $H^{-1}$ is an isomorphism as well, and thus we further have
$s\in\ \uparrow_{\mathfrak{B}}Ha\quad$ $\Leftrightarrow\quad Ha=s^{\mathfrak{B}}\mathbf{o},\quad\text{forsome $\mathbf{o}\in A^{r(s)}$},$
$\Rightarrow\quad a=H^{-1}s^{\mathfrak{B}}\mathbf{o}=s^{\mathfrak{A}}H^{-1}\mathbf{o}$
$\Rightarrow\quad s\in\ \uparrow_{\mathfrak{A}}a.$
Theorem 19 (First Isomorphism Theorem).
For any homomorphism $H:\mathfrak{A\to B}$ and element $a\in A$,
$a\lesssim_{\mathfrak{AB}}Ha.$
In case $H$ is an isomorphism, we have
$a\approx_{\mathfrak{AB}}Ha.$
Proof.
An immediate consequence of the Homomorphism Lemma 18 and (2). ∎
Example 20.
Let $\mathfrak{A}:=(\{a,b\},f)$ and $\mathfrak{B}:=(\{c\},f)$ be given by
{tikzpicture}
The mapping $H:\mathfrak{A\to B}$ is obviously a homomorphism. However, we have
$Ha\uparrow_{\mathfrak{BA}}a$ =$=\{x\}\subsetneq\left\{f^{n}x\;\middle|\;n\geq 0\right\}=Ha\uparrow_{\mathfrak{BA}}b$
which shows
$Ha\not\lesssim_{\mathfrak{BA}}a$
and thus
$a\not\approx_{\mathfrak{AB}}Ha.$
Theorem 21 (Second Isomorphism Theorem).
For any isomorphism $H:\mathfrak{A\to B}$ and elements $a,b\in A$,
$a\approx_{\mathfrak{A}}b\quad\Leftrightarrow\quad Ha\approx_{\mathfrak{B}}Hb.$
Proof.
An immediate consequence of the Homomorphism Lemma 18. ∎
6. The $(k,\ell)$-fragments
Since computing similarity is rather difficult in general, it is reasonable to study fragments of the framework. For this, we introduce in this section the $(k,\ell)$-fragments. Formally, let $X_{k}:=\{x_{1},\ldots,x_{k}\}$, for some $k,\ell\in\mathbb{N}\cup\{\infty\}$ so that $X_{\infty}=X$.
Definition 22.
Define
$\uparrow^{(k,\ell)}_{\mathfrak{A}}a:=(\uparrow_{\mathfrak{A}}a)\cap\{s(x_{1},\ldots,x_{k})\in T_{L}(X_{k})\mid\text{each of the $k$ variablesin $X_{k}$ occurs at most $\ell$ times in $s$}\}.$
We write $k$ instead of $(k,\infty)$ so that
$\uparrow^{k}_{\mathfrak{A}}a=(\uparrow_{\mathfrak{A}}a)\cap T_{L}(X_{k}).$
Every $(k,\ell)$-fragment gives raise to a similarity relation $\approx_{(k,\ell)}$ defined as $\approx$ with $X$ replaced by $X_{k}$ and $\uparrow$ replaced by $\uparrow^{(k,\ell)}$. The simplest fragment — namely, the $(1,1)$-fragment — contains only monolinear generalizations containing exactly one occurrence of a single variable $x$. We denote the set of monolinear generalizations of $a$ in $\mathfrak{A}$ by $\uparrow^{m}_{\mathfrak{A}}a$ and the so-obtained similarity relation by $\approx_{m}$.
Example 23.
In the domain of sets, we obtain the following characterization of similarity in the monolinear fragment. Let $U$ be a set and $A,B\subseteq U$. We have (here $X$ denotes a variable and not the set of variables as above)
($\uparrow^{m}_{(2^{U},\cup,2^{U})}A=\left\{X\cup B\;\middle|\;A=B\cup C,\;B,C\in 2^{U}\right\}\cup\{A\}=\left\{X\cup B\mid B\in 2^{A}\right\}\cup\{A\},$
which can be identified with the power set of $A$. Hence, for $A\neq U$, we have
[…]
$\Leftrightarrow\quad A\subseteq B,$
which implies, for $A\neq U$ and $B\neq U$,
$A\approx_{(2^{U},\cup,2^{U}),m}B\quad\Leftrightarrow\quad A\subseteq B\quad\text{and}\quad B\subseteq A\quad\Leftrightarrow\quad A=B.$
For $U$, on the other hand, we have
[…]
Example 24.
In the multiplicative algebra $\mathfrak{M}:=(\mathbb{N},\cdot,\mathbb{N})$, we have
$\uparrow^{m}_{\mathfrak{M}}a$ =$=\{kx\mid a=ko,\;o\in\mathbb{N}\}$
=$=\{kx\mid\text{$k$ divides $a$}\}.$
Hence,
$a\uparrow^{m}_{\mathfrak{M}}b=\{kx\mid k\text{ divides $a$ and $b$}\},$
which is $b$-maximal iff $a$ divides $b$, that is,
$a\lesssim_{\mathfrak{M},m}b\quad\Leftrightarrow\quad\text{$a$ divides $b$}.$
That is, within the monolinear fragment the relation $\lesssim_{m}$ captures divisibility of natural numbers in $\mathfrak{M}$.
7. Finite unary algebras aka semiautomata
In this section, we study algebraic similarity in finite unary algebras, which can be seen as semiautomata, and show that we can use well-known methods from the theory of finite automata.
In the rest of this section, let
$\mathfrak{A}=(A,\Sigma:=\{f_{1},\ldots,f_{n}\}),$
for some $n\geq 1$, be a finite unary algebra with finite universe $A$. We shall now recall that every such algebra is essentially a semiautomaton.
Recall that a (finite deterministic) semiautomaton (see e.g. ?, §2.1) is a construct
$\mathfrak{S}=(S,\Sigma,\delta),$
where $S$ is a finite set of states, $\Sigma$ is a finite input alphabet, and $\delta:S\times\Sigma\to S$ is a transition function. Every semiautomaton can be seen as a finite unary algebra in the following well-known way: every symbol $\sigma\in\Sigma$ induces a unary function $\sigma^{\mathfrak{S}}:S\to S$ via $\sigma^{\mathfrak{S}}:=\delta(x,\sigma)$. We can now omit $\delta$ and define
$\mathfrak{S}^{\prime}:=(S,\Sigma^{\mathfrak{S}}:=\{\sigma^{\mathfrak{S}}\mid\sigma\in\Sigma\}).$
It is immediate from the construction that $\mathfrak{S}$ and $\mathfrak{S}^{\prime}$ represent essentially the same semiautomaton and that every semiautomaton can be represented in that way — the difference is that $\mathfrak{S}^{\prime}$ is a finite unary algebra!
Recall that a (finite deterministic) automaton (see e.g. ?, §1.1) is a construct
$\mathscr{A}:=(Q,\Sigma,\delta,q_{0},F),$
where $(Q,\Sigma,\delta)$ is a semiautomaton, $q_{0}\in Q$ is the initial state, and $F\subseteq Q$ is a set of final states. The behavior of $\mathscr{A}$ is given by
$||\mathscr{A}||:=\{w\in\Sigma^{\ast}\mid\delta^{\ast}(q_{0},w)\in F\},$
where $\delta^{\ast}:Q\times\Sigma^{\ast}\to Q$ is defined recursively as follows, for $q\in Q,a\in\Sigma,w\in\Sigma^{\ast}$:
$\delta^{\ast}(q,\varepsilon)$ $:=q,$
$\delta^{\ast}(q,aw)$ $:=\delta^{\ast}(\delta(q,a),w).$
Notice that since automata are built from semiautomata by adding an initial state and a set of final states, and since every semiautomaton $\mathfrak{S}=(S,\Sigma,\delta)$ can be represented in the form of a finite unary algebra $\mathfrak{S}^{\prime}=(S,\Sigma^{\mathfrak{S}})$ as above, we can reformulate every automaton $\mathscr{A}=(Q,\Sigma,\delta,q_{0},F)$ as $\mathscr{A}^{\prime}=(Q,\Sigma^{\mathscr{A}},q_{0},F)$, where $\Sigma^{\mathscr{A}}:=\{\sigma^{\mathscr{A}}\mid\sigma\in\Sigma\}$ and $\sigma^{\mathscr{A}}:=\delta(\,.\,,\sigma):Q\to Q$. In other words, given a finite unary algebra (semiautomaton)
$\mathfrak{A}=(A,\Sigma),$
we can construct a finite automaton
$\mathfrak{A}_{a\to F}=(A,\Sigma,a,F)$
by designating a state $a\in A$ as the initial state, and by designating a set of states $F\subseteq A$ as final states.
We want to compute the set of generalizations $\uparrow_{\mathfrak{A}}a$. Notice that we can identify, for example, each term in $T_{\{f,g\}}(\{x\})$ with a word over the alphabet $\Sigma=\{f,g\}$: for instance, the term $fgfx$ can be identified with the word $fgf\in\Sigma^{\ast}$ since the variable $x$ contains no information. We denote the function induced by a word $w\in\Sigma^{\ast}$ in $\mathfrak{A}$ by $w^{\mathfrak{A}}$ — for example, $(fg)^{\mathfrak{A}}$ is the function on $A$ which first applies $g$ and then $f$. We shall now show that in any finite unary algebra (semiautomaton) $\mathfrak{A}=(A,\Sigma)$, the set of generalizations $\uparrow_{\mathfrak{A}}a$ can be computed by some finite automaton as illustrated by the following example:
Example 25.
Consider the finite unary algebra (semiautomaton)
$\mathfrak{A}=(\{a,b\},\Sigma:=\{f,g\})$
given by
{tikzpicture}
We can identify the set of all generalizations of $a$ in $\mathfrak{A}$ with
$\uparrow_{\mathfrak{A}}a=$$\left\{w\in\Sigma^{\ast}\;\middle|\;\delta^{\ast}(a,w)=a\right\}\cup\left\{u\in\Sigma^{\ast}\;\middle|\;\delta^{\ast}(b,u)=a\right\}.$
Now define the automaton $\mathfrak{A}_{a\to\{a\}}$ by adding to the semiautomaton $\mathfrak{A}$ the initial state $a$ and the set of final states $\{a\}$ (we use here the standard pictorial notation for automata)
{tikzpicture}
and the automaton $\mathfrak{A}_{b\to\{a\}}$ by
{tikzpicture}
We then clearly have
$\uparrow_{\mathfrak{A}}a=||\mathfrak{A}_{a\to\{a\}}||\cup||\mathfrak{A}_{b\to\{a\}}||.$
It is straightforward to generalize the construction in Example 25:
Definition 26.
Given a finite unary algebra (semiautomaton) $\mathfrak{A}=(A,\Sigma)$, the automaton $\mathfrak{A}_{b\to\{a\}}$ is the automaton induced by the functions in $\Sigma$ with start state $b$ and single final state $a$ given by
$\mathfrak{A}_{b\to\{a\}}:=(A,\Sigma,b,\{a\}).$
Fact 27.
Given any finite [^4] Finiteness is required since regular languages are not closed under infinite union. unary algebra (semiautomaton) $\mathfrak{A}=(A,\Sigma)$ and $a\in A$, we have
(6) $\uparrow_{\mathfrak{A}}a=\bigcup_{b\in A}||\mathfrak{A}_{b\to\{a\}}||.$
By Fact 27, computing
$a\uparrow_{\mathfrak{AB}}b=(\uparrow_{\mathfrak{A}}a)\cap(\uparrow_{\mathfrak{B}}b)$
amounts to computing the intersection of two regular languages by applying standard techniques from automata theory. inline]Hence, given $a\in A$ and $b\in B$, we can check
$a\stackrel{{\scriptstyle?}}{{\lesssim}}b$
by first checking … and in case … by checking
$a\uparrow b\stackrel{{\scriptstyle?}}{{\subsetneq}}a\uparrow c$
for all the finitely many $c\neq b\in B$… In total, we have thus derived an algorithm for the computation of similarity.
8. Finite algebras
In this section, we interpret algebraic similarity in finite algebras and show that it is closely related to regular tree languages and finite tree automata.
Recall that a (frontier-to-root) tree automaton $\mathfrak{T}_{k,\alpha,F}(\mathfrak{A}):=(\mathfrak{A},L,X_{k},\alpha,F)$ consists of (e.g. ?)
  • a finite $L$-algebra $\mathfrak{A}$,
  • an initial assignment $\alpha:X_{k}\to\mathfrak{A}$, and
  • a set $F\subseteq A$ of final states.
The regular tree language recognized by $\mathfrak{T}_{k,\alpha,F}(\mathfrak{A})$ is given by
$||\mathfrak{T}_{k,\alpha,F}||:=\left\{s\in T_{L}(X_{k})\;\middle|\;s^{\mathfrak{A}}\alpha\in F\right\}.$
Notice that we can write the set of $k$-generalizations of $a$ in $\mathfrak{A}$ as
$\uparrow^{k}_{\mathfrak{A}}a=\left\{s\in T_{L}(X_{k})\;\middle|\;s^{\mathfrak{A}}\alpha=a,\text{ for some initial assignment $\alpha$}\right\}.$
We thus have
$\uparrow^{k}_{\mathfrak{A}}a=\bigcup_{\alpha\in A^{X_{k}}}||\mathfrak{T}_{\alpha,\{a\}}(\mathfrak{A})||.$
Since tree automata are closed under finite union, there is some tree automaton $\mathfrak{T}_{a}(\mathfrak{A})$ such that
$\uparrow^{k}_{\mathfrak{A}}a=||\mathfrak{T}_{k,a}(\mathfrak{A})||,$
and since tree automata are closed under intersection, there is some tree automaton $\mathfrak{T}_{a,b}(\mathfrak{AB})$ such that
$a\uparrow_{\mathfrak{AB}}b=||\mathfrak{T}_{k,a,b}(\mathfrak{AB})||.$
Since the underlying algebras are finite, checking $a$-, and $b$-maximality can be done in time linear to the size of the algebra. In total, we have thus derived an algorithm for the computation of $k$-similarity in finite algebras.
9. Logical interpretation
In this section, we show how the purely algebraic notion of similarity can be naturally interpreted within the logical setting of first-order logic via the well-known concept of a model-theoretic type. For this, we associate with every $L$-term $s\mathbf{z}$ a g-formula of the form
$\varphi_{s\mathbf{z}}\,x:\equiv(\exists\mathbf{z})(x=s\mathbf{z}).$
We denote the set of all g-formulas over $L$ by $g\text{-}Fm_{L}$.
Now that we have defined g-formulas, we continue by translating sets of generalizations into sets of g-formulas via a restricted notion of the well-known model-theoretic types (e.g. ?, §7.1). Define the g-type of an element $a\in A$ by
$g\text{-}Type_{\mathfrak{A}}(a):=\left\{\varphi\in g\text{-}Fm_{L}\;\middle|\;\mathfrak{A}\models\varphi a\right\}.$
We have the following correspondence between sets of generalizations and g-types:
$s\in\ \uparrow_{\mathfrak{A}}a\quad\Leftrightarrow\quad\varphi_{s}\in g\text{-}Type_{\mathfrak{A}}(a).$
This is interesting as it shows that algebraic generalizations, which are motivated by simple examples, have an intuitive logical meaning. The expression “$a$ is similar to $b$” can thus be reinterpreted from a logical point of view by defining the relation $\approx_{T}$ in the same way as $\approx$ with $\uparrow$ replaced by $g\text{-}Type$. We then have the logical characterization of similarity in terms of model-theoretic types via
$a\approx_{\mathfrak{AB}}b\quad\Leftrightarrow\quad a\approx_{\mathfrak{AB},T}b.$
10. Related work
The idea of this paper to use generalizations to define similarity in an algebraic setting appears to be original.
The theory of anti-unification (?, ?) (cf. ?) is the theory of generalization. The name is derived from the fact that computing the generalization of two terms is the dual of computing their unification, an operation heavily studied in the literature with applications to automated theorem proving and logic programming (cf. ?). The main focus of anti-unification is the computation of (complete sets of) least or minimally general generalizations. The problem of computing or representing the set of all generalizations has, to the best of our knowledge, not been studied in anti-unification theory. In fact, Temur Kutsia — an expert in the field of anti-unification — thinks that existing tools cannot be adopted in a straightforward manner and that therefore novel tools will be needed.
There is a vast literature on similarity, most of which deals with quantitative approaches where similarity is numerically measured in one way or another. A notable exception is ? (?) where qualitative judgements of similarity are interpreted through statements of the form “$a$ is more similar to $b$ than $c$ to $d$.” ? (?) analyze the role of qualitative similarity in analogical transfer.
11. Future work
At the core of similarity is the set $\uparrow_{\mathfrak{A}}a$ of all generalizations of an element $a$ in an algebra $\mathfrak{A}$. It appears challenging to compute $\uparrow_{\mathfrak{A}}a$ even in simple concrete domains of infinite cardinality like the multiplicative algebra or the word domain. A reasonable starting point are unary algebras for which we have established, in the finite case, a close relationship to regular languages and finite automata in §7. Computing sets of characteristic generalizations is equally important since the set of all generalizations is in most cases infinite and complicated to compute.
We have seen in Example 20 that the First and Second Isomorphism Theorems 19, 21 can in general not be generalized to homomorphisms. Are there algebras in which homomorphisms are compatible with similarity and how can they be characterized? More generally speaking, what kind of mappings are compatible with similarity?
Another possibility is to study fragments of the framework where generalizations are syntactically restricted. In §6, we sketched the monolinear fragment where generalizations contain only a single occurrence of a single variable. The linear fragment consisting of generalizations containing at most one occurrence of multiple variables is the next logical step.
A practically interesting domain of study relevant for applications in computational linguistics, biology, and AI in general, is the word domain. Again, computing the set of all (characteristic) generalizations of a given word appears challenging. Nonetheless, it is interesting to study similarity of words and how it is related to the numerous notions of word similarity (and distance) in the literature (cf. ?).
We have seen in Theorem 16 that similarity is in general not transitive justified by a simple counterexample. In those algebras where similarity is transitive, it forms an equivalence relation and enjoys all the nice properties that come with it. It is therefore desirable to have general criteria under which algebras induce a transitive similarity relation.
Finally, recall that the two algebras $\mathfrak{A}$ and $\mathfrak{B}$ are assumed to have the same underlying language. This is essential for the computation of $(\uparrow_{\mathfrak{A}}a)\cap(\uparrow_{\mathfrak{B}}b)$, which is at the core of the framework. Generalizing the framework to two languages $L_{\mathfrak{A}}\neq L_{\mathfrak{B}}$ appears challenging.
12. Conclusion
This paper introduced from first principles an abstract algebraic and qualitative notion of similarity within the general setting of universal algebra based on sets of generalizations motivated by the observation that these sets encode important properties of elements. It turned out that similarity defined in this way has appealing mathematical properties. We showed that similarity is compatible with structure-preserving mappings; can be computed in finite algebras using tree automata; and we gave a logical interpretation of similarity in terms of model-theoretic types.